OpenStudy (nicole143):

If you are trying to find all the zeros of an equation and your last "integer" doesn't have an "X" do you just leave it has is and not simple before putting it into the quadratic equation? @mathmale

OpenStudy (mathmale):

Here's my interpretation of what you're asking: "If you're given a polynomial equation in which the last COEFFICIENT is a constant (that is, it doesn't stand before any power of x), do you leave that CONSTANT COEFFICIENT as is?" Actually, Nicole, if I'm right on this interpretation, this is the situation you'll almost always encounter.

OpenStudy (mathmale):

For example: Find all the zeros of the polynomial equation x^2 - x - 6 = 0. You'd definitely want to leave the equation as is. What methods of factoring quadatric equations are you familiar with ?

OpenStudy (nicole143):

This being an example: After the first arrow it is simplified cause the last "integer" has an x.

OpenStudy (nicole143):

So, would it be the same or do I have to do it differently since I don't have an "X" on the last bit?

OpenStudy (mathmale):

Oh, it helps so much to see what you've already done. Did you mean y - x^3 - 3x^2 - 9x or y = x^3 - 3x^2 -ax?

OpenStudy (nicole143):

I didn't want to put up the equation I'm working with because I wanted to do it on my own but if you need to see it I can send it to you. The picture is of an equation I have already done - the one I am working with now doesn't have an ending "X" so I can't factor it out into say.. x(x^3 + 7x^2 and continued because if I did and then you times it, it would then have an "X" that wasn't there before..

OpenStudy (mathmale):

May I refer to the most recent photo that you've shared with me? is that the problem you're working with? Just one look at that photo and I pretty much knew how to respond.

OpenStudy (mathmale):

In the long run I'd prefer that you just post on OpenStudy or send me privately photos of any problem(s) you'd like to discuss. You are to be commended for working on these problems on your own and having something to share because of your having done so.

OpenStudy (nicole143):

The equation I'm working with now is: -3x^4 + 27x^2 + 1200 = 0 I am trying to find all the zeros for the equation but because there is not an "X" on the end of 1200. I also would like to know if I have to put in place holders such as x^3?

OpenStudy (mathmale):

Thanks for typing this out. I see what you're asking. Yes, you do have to put in place holders for any missiong powers of x. Represent that missing x^3 with the term 0*x^3 and the missing x with 0*x. This fourth order polynomial should theoretically have FIVE terms on the left side.

OpenStudy (mathmale):

Are you familiar with synthetic division? If not, what method were you planning to try to find the zeros of this equation?

OpenStudy (mathmale):

Also, leave that 1200 on the end exactly as it is. If you like, you could think of that 1200 as the coefficient of x^0. Note that x^0 = 1.

OpenStudy (nicole143):

Okay, and yes I know synthetic division. I had used the equation x = -b +- the square root of b^2 - 4ac / 2a to solve the last one I did but I was factored out

OpenStudy (mathmale):

Hold just a moment...I'm thinking.

OpenStudy (mathmale):

First, please share with me the approach you were thinking of using. I don't want to push my own approaches on you if you already have one of your own that would work.

OpenStudy (nicole143):

I'm not entirely sure what to use:( Would this be right for the next step?

OpenStudy (mathmale):

Yes, if you'd please write the coefficient ZERO in front of both the x^3 and the x^1.

OpenStudy (mathmale):

Those are place holders, remember?

OpenStudy (nicole143):

Okay, I wrote them.

OpenStudy (mathmale):

May I ask where this problem came from? I see a way to solve it, but it's a challenging problem.

OpenStudy (nicole143):

I have to turn in work for school. I hadn't done one like this before and I wanted to make sure I know how and that I did it right before sending it i for scoring. If you would like - I wouldn't mind - changing the number so I would still do the work that has to be turned in on my own.

OpenStudy (mathmale):

You're surely very honest!

OpenStudy (nicole143):

So if we did the same type of problem so I learned how, I could then do the scoring one on my own. And thank you ha

OpenStudy (mathmale):

please excuse me, just a moment.

OpenStudy (nicole143):

Okay. Again, thank you for all your help!

OpenStudy (mathmale):

I believe I could help you with a similar problem, but must warn you that it'd take a while. I'd be OK, personally, discussing what to do with this particular problem, with you doing the work and I providing suggestions and guidance.

OpenStudy (mathmale):

I'll leave it up to you what we do next, if anything. My first suggestion would be that you take the original equation and divide it through by -3. Can you see any value in doing tht?

OpenStudy (nicole143):

Okay, that would be great because I also have other work to do and would like to go to bed not to late. Dividing by -3 would make the 27 a -9 and the 1200 a -400 - making the equation smaller.

OpenStudy (mathmale):

By the way, my approach worked: one solution / zero is 5. OK, Nicole, perhaps we'll have the opp to work together again later; I hope so. Admire your perseverance and your sharing your previously completed work with me. Yes, if you divide that equation through

OpenStudy (mathmale):

by -3, you'd get x^4 -27x^2 - 400 = 0, which is significantly easier to solve.

OpenStudy (nicole143):

If you have to go, could you please quick tell me what to do next and I'll hope I don't mess up on making my way to the end.

OpenStudy (mathmale):

More power to you for being brave enought o do that . I'm thinking, just for a moment, about what to tell you next.

OpenStudy (mathmale):

Nicole: try again dividing the original equation by -3. That -27 is incorrect.

OpenStudy (nicole143):

27 / -3 = -9 though - I did it on my calculator because I'm not good with negative ha

OpenStudy (mathmale):

OK. So, are we in agreement that we have x^4 - 9x^2 - 400 = 0?

OpenStudy (nicole143):

Yes, that is what I got after / by -3

OpenStudy (mathmale):

Cool. Now you're exactly on the right track. This latest equation is easier to understand and to solve than the original one.

OpenStudy (mathmale):

OK: let me ask how comfortable you would be in letting p = x^2, so that p^2 - 9p - 400 = 0. Remember having done that before?

OpenStudy (nicole143):

I didn't do that before, I hadn't see why to since you would just put it back later. I may be mistaken though.

OpenStudy (nicole143):

I haven't ever used that

OpenStudy (mathmale):

It's a commonly used trick. Which looks easier to solve, x^4 - 9 x^2 - 400 = 0, or p^2 - 9p - 400 = 0?

OpenStudy (nicole143):

To me; x^4 but only because I haven't used P in this certain type of equation but if it makes it easier, then I'll try.

OpenStudy (mathmale):

In other words, making the substitution p = x^2 reduces the fourth order polynomial to a 2nd order one (a quadratic), which I'm sure you'd be able to solve. Suppose you were to get p = 25 (which actually is a solution); then p = 25 = x^2 and x would then be 5 or -5. Are you comfortable with this reasoning?

OpenStudy (mathmale):

So, p = 25 is just an intermediate result; We'd have to remember that p = x^2 and that 25 = x^2, so that x = plus or minus 5.

OpenStudy (nicole143):

I think I get it.

OpenStudy (nicole143):

I'm not sure what to do have you put p = x^2 in

OpenStudy (mathmale):

I believe there are two more soltions, but could be wrong on that. Again, the reason I make that substitution is to reduce the order of the original polynomial from 4 to 2. You do NOT have to do it this way. You could use synthetic division; Try it: Let 5 b e

OpenStudy (nicole143):

Do you add 400 to both sides?

OpenStudy (mathmale):

No. We want to keep that equation in standard form. Adding 400 to both sides would mess up that standard form.

OpenStudy (nicole143):

Oh, I see.

OpenStudy (mathmale):

If you're comfortable with synthetic division, please go ahead and try verifying that 5 is a root of the original equation. 5 | 1 0 -9 -400

OpenStudy (mathmale):

Want to give it a try?

OpenStudy (nicole143):

Yes

OpenStudy (mathmale):

I've just done that, and get a zero remainder, which verifies that 5 is indeed a root of the original equation.

OpenStudy (mathmale):

Do you get the same result thru synthetic division?

OpenStudy (nicole143):

I think I did it wrong ha :( I got -320, NOT good with negatives..

OpenStudy (mathmale):

Want to give it another try? Practice makes perfect.

OpenStudy (nicole143):

Yes ha

OpenStudy (mathmale):

5 | 1 0 -9 0 -400 results in what remainder?

OpenStudy (nicole143):

One Second, I think I missed a zero

OpenStudy (mathmale):

5 | 1 0 -9 0 -400 (for better visibility)

OpenStudy (nicole143):

Do you add or subtract? Just to be clear

OpenStudy (mathmale):

Depends on the sign of each particular coefficient {1, 0, -9, 0, -400}.

OpenStudy (nicole143):

I think thats where I am going wrong..

OpenStudy (mathmale):

Bring that 1 down. Multiply that 1 by 5. Result is positive, right? write that 5 under the first 0. Add 0 and 5: you get 5. Now mult that 5 by 5 and write the result under the -9. In that case you'd have 25-9 = 16. Try finishing the problem.

OpenStudy (nicole143):

OKay

OpenStudy (mathmale):

5 times 16 is 80; write the 80 under the 2nd zero coefficient and add the two numbers together. Result: 80. mult that 80 by 5 and write the result under the -400. OK?

OpenStudy (nicole143):

:D I got it! I had done it like that but was missing the 2nd zero

OpenStudy (mathmale):

So cool! Bully for you, Nicole!! Now you should see a new string of coefficients : 1 5 16 80. Do you?

OpenStudy (nicole143):

Yes

OpenStudy (mathmale):

Great. Perfect.

OpenStudy (mathmale):

Now test my thesis that -5 is also a root of the original equation

OpenStudy (nicole143):

Pu tin -5 as "e"?

OpenStudy (nicole143):

*put

OpenStudy (mathmale):

by performing the following synthetic division: -5 | 1 5 16 80. Go ahead, please.

OpenStudy (nicole143):

Okay

OpenStudy (nicole143):

I end up with a 0 under the -80

OpenStudy (mathmale):

Fantastic. What does that mean?

OpenStudy (nicole143):

That both 5 and -5 are zeros for the equation

OpenStudy (mathmale):

Absolutely right! Congratulations!!!

OpenStudy (mathmale):

Now you have one final row of coeffciients: they are 1 0 16, right?

OpenStudy (nicole143):

Yes

OpenStudy (mathmale):

So, what does that mean for us? 1 0 16 translates into the quadratic 1x^2 + 16 = 0, which has only imaginary roots. If you were supposed to find ALL of the roots, then you'd need to write the results as {5, -5, 4i, -4i} and you'd be done!

OpenStudy (mathmale):

You could, if you wished, verify through synth div that all four of these numerals are indeed roots of the original equation!

OpenStudy (nicole143):

Great! Thank you soo much for all your help! I'll will put it all together and hope I do as well on the rest as I did for this one(:

OpenStudy (mathmale):

My great pleasure. Really! You're so cool to work with....quick, willing to work, able to express yourself very well and to find your own mistakes.

OpenStudy (mathmale):

Post any further questions you have right here in OpenStudy; perhaps I'll have the good fortunate to see them and to work with you further. Happy New Year to you, Nicole, and good night!

OpenStudy (nicole143):

Happy New Years to you as well! Have a goodnight.

OpenStudy (mathmale):

Bye!